# Absence of bulk charge density wave order in the normal state of UTe2

### v2

@@ Acknowledgments
WS acknowledges fruitful discussions with Aline Ramirez. WS was supported through funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 884104 (PSI-FELLOW-III-3i). Work at Los Alamos National Laboratory was performed under the auspices of the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Science and Engineering. C.S.K. acknowledges support from the Laboratory Directed Research and Development program. P.A. acknowledges support from the Gordon and Betty Moore Foundation, EPiQS grant GBMF9452. J. C. acknowledges support from Swiss National Science foundation under grant 200021-188564. MJ acknowledges funding by the Swiss National Science Foundation through the project ”Berry-Phase Tuning in Heavy f-Electron Metals” (200650). We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III at DESY. Beamtime was allocated for proposal I-20221340 EC.
Author information
The authors declare no competing financial interests. Correspondence and requests for materials should be addressed to ckengle@lanl.gov and wsimeth@lanl.gov.

### v3

@@ Acknowledgments
WS acknowledges fruitful discussions with Aline Ramirez. The authors further thank Simon Gerber for discussions. WS was supported through funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 884104 (PSI-FELLOW-III-3i). Work at Los Alamos National Laboratory was performed under the auspices of the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Science and Engineering. C.S.K. acknowledges support from the Laboratory Directed Research and Development program. P.A. acknowledges support from the Gordon and Betty Moore Foundation, EPiQS grant GBMF9452. J. C. acknowledges support from Swiss National Science foundation under grant 200021-188564. MJ acknowledges funding by the Swiss National Science Foundation through the project ”Berry-Phase Tuning in Heavy f-Electron Metals” (200650). We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III at DESY. Beamtime was allocated for proposal I-20221340 EC. Author contributions: MJ, PFSR, and WS conceived the study. MJ and WS designed the experiments. PFSR grew the samples. CK, JV, SF, and WS carried out the X-ray experiments. CK, PFSR and WS interpreted the results. CK, PFSR, and WS wrote the paper with input from all the authors. Competing interests: The authors declare that they have no competing interests. Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Correspondence and requests for additional materials may be addressed to wsimeth@lanl.gov.